an amplifier's open circuit output voltage is 10V peak to peak. if when a 10KΩ load is connected the output voltage drops to 8V peak to peak determine the amplifier's output resistance
1
Expert's answer
2020-10-15T10:50:35-0400

Since the voltage dropped from "10 V" to "8 V" , therefore, the voltage "2 V" at the new load, therefore the current at the load "I =\frac {U _ {R}} {R} =\frac {2} {10 ^ 4} = 0.0002" A.

Therefore, resistor resistance "R_{amplifier}=\frac{U_{output}}{I}=\frac{8}{0.0002}=40" K"\Omega".